A spacecraft with the shape of a sphere of diameter D moves past an observer on Earth with a speed of 0.5c. What shape does the observer measure for the spacecraft as it moves past?
1
Expert's answer
2015-05-13T02:49:51-0400

Solution:
L'=L* sqrt(1- (v/c)^2)
If v=0.5c - > L'= L*sqrt(1- (0.5c/c)^2) = L * 0.866
L = D -> L' = 0.866*D

So the shape does the observer measure for the spacecraft as it moves past is ellipse with major axis D and minor axis 0.866D